OpenStudy (anonymous):

help How can x2 = x2 + 2x + 9 be set up as a system of equations? A) y = x2 − 9 y = x2 + 2x + 9 B) y = x2 y = x2 + 2x + 9 C) y = x2 + 2x y = x2 + 9 D) y = x2 y = 2x + 9

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

x2 means x^2

OpenStudy (zzr0ck3r):

Any ideas?

OpenStudy (anonymous):

no not at all

OpenStudy (anonymous):

I was thinking option B but I'm not sure

OpenStudy (zzr0ck3r):

The way you can check, is by setting each equation in the system equal to each other and see if you get the above equation Example: The first one y = x2 − 9 y = x2 + 2x + 9 Set them equal x^2-9=x^2+2x+9 This is not what we started with, so it is not the answer.

OpenStudy (zzr0ck3r):

You are right, but now do you understand why?

OpenStudy (anonymous):

yes I do :)

OpenStudy (anonymous):

so letter B is the correct answer

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

if you set them equal, you get the thing you started with.